A taxi charge a fixed amount as rental for the first 3 km and it charges a fixed amount for every additional km. A passenger paid Rs. 31.40 for travelling 25 km. And another passenger travelled 30 km paid Rs 37.40. How will a passenger have to pay for travelling 55 km?
Answer by macston(5194)

F=fixed fee (first 3 km); R=rate per kilometer after 3km
F+(total mileage - 3km)(R)=Total fare
F+22R=31.40
F+27R=37.40 Subtract these two equations
-5R=-6.00
R=Rs. 1.20 The rate for over 3 km is Rs. 1.20/km
F+22R=Rs. 31.40
F+22(Rs. 1.20)=Rs. 31.40
F+Rs. 26.40=Rs. 31.40
F=Rs. 5.00 The rate for the first 3 km is Rs. 5.00
.
For total mileage=55 km:
Total fare=F+(mileage-3km)R
Total fare=Rs. 5.00+(55km-3km)(Rs. 1.20/km)
Total fare=Rs. 5.00 + Rs. 62.40=Rs. 67.40
ANSWER: The fare for a 55 km trip would be Rs. 67.40.